OpenStudy (anonymous):

A segment with endpoints C (3, 4) and D (11, 3) is divided by a point E such that CE and DE form a 3:5 ratio. Find the x value for E. 6 5.5 7.5 7

OpenStudy (anonymous):

@waterineyes @dan815 @abb0t

OpenStudy (abb0t):

@dan815

OpenStudy (anonymous):

@abb0t can you help please

OpenStudy (anonymous):

@ash2326 @ironictoaster @amistre64 @annas

OpenStudy (anonymous):

@nikato please

OpenStudy (amistre64):

move it to the origin to help measure it ...., then divide it into the requisite parts

OpenStudy (anonymous):

i got 7

OpenStudy (amistre64):

C (3, 4) and D (11, 3) -3 -4 -3 -4 ------------------- 0 0 8 -1 now, we see that the x parts are from 0 to 8, to it is 8 units long. a ratio of 3:5 IS 8 parts, 3 to the left and 5 to the right let Ex = 3, and move it back into place 3, y +3 +4 ------- 6, y+4 is what E would look like given whatever the y part wants to be

OpenStudy (anonymous):

oh so its 6? okay thanks @amistre64

OpenStudy (anonymous):

Write an equation of a line parallel to y = 1 over 4x − 6 in slope-intercept form that passes through the point (−1, 5). y= 1 over 4x − 1 y= 1 over 4x + 5.25 y= −4x + 5 y= −4x − 1.25

OpenStudy (amistre64):

please ask new question in a new post ....